OpenStudy (anonymous):

Solve x2 − 16 = 0 and x2 = −2x + 24. Show your work?

OpenStudy (anonymous):

\[x^2-16=0\] \[x^2=-2x+24\] Use substitution. Or do something cute: \[x^2-16=(x+4)(x-4)=0\] Follow this from here.